Please help me with this question:
A water main for a street is being laid using a particular kind of pipe that comes in either 18ft or 20ft sections. The designer has determined that the water main would require 14 fewer sections of 20 ft pipe than if the 18 ft sections were used. Find the total length.
At first I thought maybe it was a variable for each section like 18x=20x-14x... but that doesn't work
then I tired with 2 variables of 18x=20y-14y... I ended up with y=3x and when I subbed that in I got x=1/3. but its not right. can you help me find the correct equation to use for this problem?
Thank you

Answer by stanbon(75887)

A water main for a street is being laid using a particular kind of pipe that comes in either 18ft or 20ft sections.
--------------------------------------------------
The designer has determined that the water main would require 14 fewer sections of 20 ft pipe than if the 18 ft sections were used.
-----
Find the total length.
---
Let the total length be x ft.
----
# of 20 ft sections would be x/20
# of 18 ft sections would be x/18
----
Equation:
# of 18 - # of 20 = 14
x/18 - x/20 = 14
---
Multiply thru by 18*20 to get:
------
20x - 18x = 14*18*20
2x = 14*18*20
----
x = 2520 ft (total length)
==============================
Cheers,
Stan H.